1.         For each set below, using Excel, construct

           a.         a frequency distribution,

   b.         a relative frequency distribution, and

           c.         a cumulative relative frequency distribution.

Consider whether or not you should group your data. Describe how you determined your bin width, if you grouped the data in intervals.

Set 1

75   

95   

103

100

93

91

90

92

89

105

86

85

81

96

103

99

94

95

91

97

92

107

75

106

92

97

88

87

89

104

97

113

93

108

101

84

96

83

98

115

109

98

88

110

99

94

103

79

95

114

118

102

104

Set 2

82

101

89

114

128

118

105

91

100

99

113

106

106

102

95

120

97

106

103

105

112

98

85

93

107

119

104

110

97

109

100

109

95

115

83

103

90

108

83


Set 3

0   

2

0

1

0

0

0

0

1

0

0

0

0

0

0

2

0

0

0

0

2

0

1

0

0

0

4

1

0

0

1

0

0

0

1

0


2. Assume that the first set of scores below represents the IQ scores of children

who have been participating in a school lunch program and the second set of scores represent the IQs of a similar group of children who have not participated in the program.

a.         Prepare appropriate graphical representations of the frequency and relative frequency distributions.

b.         Compare the two distributions by placing them on the same graph. Be sure to have equivalent intervals, if grouped. Why did you decide to compare frequencies, or relative frequencies?

3.         Assume that the third set below are data from a survey of riders of a city bus system. Each rider was asked to indicate the number of times he or she had taken a bus on a Sunday during the previous month.   Prepare a graphical representation of the relative frequency data.
Answer #1

(1)

Set 1

Class size = (Max - Min)/Number of classes = (118 - 75)/5 = 8.6 (say 10)

CI F Relative F Cumulative Relative F
75 - 85 6 0.1132 0.1132
85 - 95 17 0.3208 0.4340
95 - 105 20 0.3774 0.8113
105 - 115 8 0.1509 0.9623
115 - 125 2 0.0377 1.0000
Sums = 53 1

Set 2

Class size = (Max - Min)/Number of classes = (128 - 82)/5 = 9.2 (say 10)

CI F Relative F Cumulative Relative F
80 - 90 5 0.1282 0.1282
90 - 100 9 0.2308 0.3590
100 - 110 16 0.4103 0.7692
110 - 120 7 0.1795 0.9487
120 - 130 2 0.0513 1.0000
Sums = 39 1

Set 3

Here there is no need to make class intervals

X F Relative F Cumulative Relative F
0 26 0.7222 0.7222
1 6 0.1667 0.8889
2 3 0.0833 0.9722
4 1 0.0278 1.0000
Sums = 36 1